How many ounces of a 22% alcohol solution mixed with how many ounces of pure alcohol would form 60 ounces of a 35% solution?
Answer by Paul(988)

22(x+60)=35(60)
22x+1320=2100
22x=750
x=34

Hence, about 34 oncues of pure alcohol should be mixed.
Paul.